# references.bib

### v1

@@ 3.   The Gated Perceptron for Computing Linear and non Linear Regression
Figure 5 presents the results of the nonlinear regression computation using the two variables $(x_{1},x_{2})$ corresponding to the third and fourth columns of the Iris dataset [[5]]. The boundary is determined with high accuracy. Three out of fifty elements from the class "Iris-versicolor" are misclassified, and only one out of fifty elements from the class "Iris-virginica" is misclassified. All elements of the class "Iris-setosa" are correctly classified. The learning rate $(lr)$ is set to $(0.05)$ over $40$ epochs.
@@ 4.1.2 Discussion
The obtained results compete those obtained in published methods so far. We can cite the most relevant methods and scores in the following table.
In [[12]] the classifiers Support vector machine (SVM), Random Forest (RF), K-nearest neighbors(K-NN), Decision tree (DT), Naïve Bayes (NB), Logistic Regression (LR), AdaBoost (AB), Gradient Boosting (GB), Multi-layer perceptron (MLP), Nearest Cluster Classifier (NCC), and voting classifier (VC) have been used for comparing and analyzing breast cancer into benign and malignant tumors. the result shows that the Voting classifier has the highest accuracy, which is 98.77%, with the lowest error rate. The results are given by table 3.

### v2

@@ 4.1.2 Discussion
The obtained results compete those obtained in published methods so far. We can find the most relevant methods and scores in [[12]]. The classifiers Support vector machine (SVM), Random Forest (RF), K-nearest neighbors(K-NN), Decision tree (DT), Naïve Bayes (NB), Logistic Regression (LR), AdaBoost (AB), Gradient Boosting (GB), Multi-layer perceptron (MLP), Nearest Cluster Classifier (NCC), and voting classifier (VC) have been used for comparing and analyzing breast cancer into benign and malignant tumors. the result shows that the Voting classifier has the highest accuracy, which is 98.77%, with the lowest error rate. The results are given by table 3.